OpenStudy (anonymous):

solve for t. S=5r^2t how do i do i solve this?

OpenStudy (anonymous):

is t part of the exponent ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@wendylisette

OpenStudy (anonymous):

what are you solving for?

OpenStudy (anonymous):

ok you have to get rid of every variable and leave t alone by doing the inverse of everything

OpenStudy (anonymous):

t = No solution

OpenStudy (anonymous):

im solving for t. these are the answer choices i have. t=25r/s t=s/5-r t=s/5r^2 t=r^2-5s

OpenStudy (anonymous):

i got the first answer for my answer but its wrong, so i am stuck on how to solve this

OpenStudy (anonymous):

so first get rid of the exponent , since it is being squared the inverse is square root. what you do to one side you MUST do to the other. therefore \[\sqrt{S} = 5rt\] because we just got rid of the exponent 2. now you have to get t alone. 5rt are all being multiplied together and the inverse is division. divide 5rt by 5r to get t alone. Remember you must do it to both sides. \[\frac{ \sqrt{S} }{ 5r } = r\] now you have r by itself

OpenStudy (anonymous):

omg your answer choices dont match ugh. lemme see where i went wrong

OpenStudy (anonymous):

ok i see , they didnt square root it

OpenStudy (anonymous):

t ∈ Ø

OpenStudy (anonymous):

it will be t = s / 5r^2

OpenStudy (isaiah.feynman):